a rope is cut in half and 1/2 is discarded. From the remaining half, 1/4 is cut off and discarded. What fraction of the original rope is left?
thanks

Answer by sudhanshu_kmr(1152)


fraction of the rope = (1 - 1/2) (1- 1/4)
= 1/2 * 3/4
= 3/8

so, 3/8 of the original.